OpenStudy (anonymous):

Help!!!! What is the sum of the geometric sequence 1, -6, 36, ... if there are 7 terms?

OpenStudy (anonymous):

Help???

OpenStudy (anonymous):

sn = a1(1 - r^n) / (1 - r) n = 7 (your looking for total of sum of numbers to the 7th term) a1 = first term = 1 r = common ratio = -6 now lets sub our info in the formula... s7 = 1(1 - (-6^7) / (1 - (-6) s7 = 1(1 - (- 279936) / (1 + 6) s7 = 1(1 + 279936) / 7 s7 = 1(279937)/7 s7 = 279937/7 s7 = 39991 <-- your answer